# CafkNet: GNN-Empowered Forward Kinematic Modeling for Cable-Driven Parallel Robots

### v2

@@ Abstract
The Cable-Driven Parallel Robots (CDPRs) have gained significant attention due to their high payload capacity and large workspace. When deploying CDPRs in practice, one of the challenges is kinematic modeling. Unlike serial mechanisms, CDPRs have a simple inverse kinematics problem but a complex forward kinematics (FK) issue. Therefore, the development of accurate and efficient FK solvers has been a prominent research focus in CDPR applications. By observing the topology within CDPRs, in this paper, we propose a graph-based representation to model CDPRs and introduce CafkNet, a fast and general FK solver, leveraging Graph Neural Network (GNN). CafkNet is extensively tested on 3D and 2D CDPRs in different configurations, both in simulators and real scenarios. The results demonstrate its ability to learn CDPRs’ internal topology and accurately solve the FK problem. Then, the zero-shot generalization from one configuration to another is validated. Also, the sim2real gap can be bridged by CafkNet using both simulation and real-world data. To the best of our knowledge, it is the first study that employs the GNN to solve FK problem for CDPRs. Videos and codes are available at https://sites.google.com/view/cafknet/site.
@@ I Introduction
Cable-driven parallel robots (CDPRs) are a kind of parallel mechanism that employs flexible cables to replace rigid links [1, 2, 3, 4]. It has exhibited excellent scalability in terms of size, payload, and dynamics capacities.
The most crucial step before deploying CDPRs is modeling, including kinematic modeling (KM) and dynamics modeling (DM). KM refers to the process of creating a mathematical equation of the end-effector movement, including its kinematic parameters such as position, orientation, and velocity. DM, on the other hand, is the process of creating a mathematical representation of how the end-effector will respond to external forces over time. For the KM of CDPR, the usual practice, like [5], is to first determine the configuration of the CDPR according to the application requirements, including the number of cables, positions of the pulley on the framework, the shape of the end-effector and the position on which the cable is connected. Then, according to the geometric relationship, the modeling of cable attachment on the end-effector and anchor points on the framework is formulated. After that, some downstream tasks can be completed, such as workspace determination [6], path planning [7], and control [8]. In CDPR’s KM, there are typically two types of calculations: forward kinematics (FK) and inverse kinematics (IK). FK involves determining the end-effector pose (i.e., position and orientation) based on cable lengths, while IK solves for the length of cables given the end-effector pose in reverse. In contrast to the traditional serial mechanical structure, the FK problem of CDPRs is more complex than its IK.
Therefore, numerous works focus on solving FK problems of CDPRs. [9] discusses the non-trivial computation of the FK of cable-driven parallel robots, which are kinematically over-constrained mechanisms. In addition, [1] provides the overview of FK solvers based on successful application cases of CDPRs. Recent work [10] proposes a neural network approach for solving the FK of under-constrained cable-driven parallel robots. In summary, solving the FK for CDPRs presents several challenges.
@@
It is obvious that due to the nonlinearity and complexity of FK problems, FK solvers have a large computational burden and also suffer from the problem of no solution or multiple solutions. Furthermore, configuration-dependent FK solvers cannot be easily applied to the reconfiguration CDPR (RCDPR) [11] or the case of simply adding or subtracting some cables. However, according to our observations in [12], we notice that the change of CDPR’s configuration does not affect the topological architecture if its configuration can be represented as a graph. In addition, it is an indisputable fact that the Sim2Real gap is inevitable, despite significant efforts to reduce differences [13, 14, 15]. This is due to the difficulty of measuring and modeling systematic errors (e.g., cable attachment position errors) and random errors (e.g., cable sagging) in CDPR, some of which are even unknowable. One promising approach currently being explored is using real data to learn about these errors, as it has been shown to be an effective way to bridge the gap.
In this paper, we propose a forward kinematics solver based on graph neural Networks (GNN) for Cable-driven parallel robots, denoted by CafkNet. By utilizing the proposed graph representation of CDPR, the CafkNet can rapidly learn the underlying topological structure of CDPR and serve as an FK solver to solve FK problems. Thanks to the topological similarity of graphs, CafkNet trained from a specific configuration can be easily transferred to other configurations without fine-tuning. Additionally, CafkNet performs well in bridging the Sim2Real gap between vast easily accessible simulation data and limited expensive real-world data. We have conducted extensive simulations and real-world experiments to demonstrate the modeling capabilities of CafkNet. To the best of our knowledge, it is the first study to employ the GNN for FK problem of CDPRs.
@@
  • We propose the CafkNet, a fast and general FK solver for CDPRs with any configurations based on GNN.
@@
The overview of CafkNet is exhibited in Fig. 1. The rest of the paper is organized as follows. Sec. II reviews related studies about KM of CDPR and GNN applications in robotics. Sec. III details the CDPR geometric model. Sec. IV introduces the proposed graph representation of CDPR and the architecture of CafkNet. Then, Sec. V demonstrates the experiment details followed by modeling results of CafkNet in Sec. VI. Finally, the conclusion and outlook are presented in Sec. VII.
Figure 1: Overview of CafkNet. (a) The geometric model of a spatial CDPR with $m$ cables. (b) The proposed graph presentation for the CDPR (see Sec. IV-A). (c) The top-down computation scheme in each message propagation block. The cable nodes collect data from the world node and are eventually aggregated by the body node (see Sec. IV-B).
@@ II-A Kinematic Modeling of CDPR
KM is a fundamental problem in CDPR’s studies and also the first step in each CDPR’s application, as revealed in [1, 2, 3]. In CDPR, the IK problem, i.e., the rope length determination given the end-effector pose, is easy to solve, relative to the FK issue. Due to the nonlinearities in FK, there are three potential situations when solving FK [16], i.e., no solution, one solution, and multiple solutions. In addition, [9] introduces an FK solver based on energy minimization for a numerical optimization technique. Also, [17] takes into account the nonlinearities of cables and proposes a method for solving the FK of CDPRs with sagging cables. However, these methods are mainly used to model CDPR more finely under their respective configurations, so as to complete the optimization of FK. More generally, some studies [18, 17] provide the FK solver using graph and graph theory to model the CDPR. Recent works employ diverse machine learning methods trying to solve the system modeling problem, such as neural networks [19, 10], deep learning [20], etc. Although their results validate the feasibility of neural networks on FK, they still need to be retrained when changing their configurations. In this work, we combine the graph and neural network and utilize GNN to represent and model the CDPR kinematics. Benefitting from GNN, the proposed CafkNet method could transfer its learned topological structure of CDPR to other configurations without retraining.
@@ II-B Graph Neural Networks in Robotics
Recently, data-driven methods have emerged as an effective approach to approximate robotics modeling. Given the fact that the robot’s structure can be easily represented as a graph, where joints are denoted as nodes and links are denoted as edges, GNNs have been increasingly applied in robotics sensing [21, 22], modeling [23, 24, 25], and control [26, 27, 28, 29].  [21, 22] propose to use GNN to encode the distributed tactile sensing information for downstream tasks. [23, 24] propose using GNN to model system dynamics, where graphs represent the rigid bodies.  [25] aims to find the representation of robot data, including kinematic data and motion data using the GNN message computation scheme. NerveNet [26] seeks to train a generalized control policy that can transferred to other robot configurations.  [27, 28, 29] propose a similar idea to train a structure-agnostic locomotion policy with a novel message-passing scheme. This paper is the first work to employ the GNN technique to solve the FK issues for the CDPRs.
III Background
@@ III-A Geometric Model of CDPR
A 6-degree-of-freedom (DoF) spatial CDPR is illustrated in Fig. 1-(a), where two endpoints of the $i$-th cable ($i=1,2,\cdots,m$) are denoted by $A_{i}$ and $B_{i}$. Given the geometric relationship, the cable vector $\mathbf{l}_{i}$ can be determined as follows.
@@ III-B Kinematics of CDPR
IV Methodology
IV-A Graph Representation of CDPR
@@
IV-B GNN-Empowered Forward Kinematic Modeling
@@
IV-C Transfer Learning across Configurations
CafkNet employs parameter sharing among identical node relationships, enhancing its ability to perform structural transfer tasks effectively. Specifically, there are only two parameterized MLPs in our graph for message propagation, including $\gamma_{wc}$ for the relationship between world node and cable node, as well as $\gamma_{cb}$ for the relationship between cable node and body node. The parameters of MLPs are independent of cable number. As a result, our approach facilitates the application of the kinematic model to various CDPR configurations, significantly improving generalization capabilities across different cable arrangements. Extensive experiments in Sec. VI verify the transferability of our CafkNet.
V Experiments
V-A Dataset
@@ V-C Implementation Details
VI Results and Discussion
VI-A FK Modeling for CDPR
@@
VI-B Transfer Learning in CDPR
@@ VI-C Bridging Sim2Real Gap
VI-C1 Simulated Data with Artificial Noises
@@
VI-C2 Real Experimental Data
@@
VII Conclusion
This paper presents the CafkNet, a GNN-empowered FK solver for CDPRs. It is the first work to utilize GNN for FK problem-solving in CDPRs. The study begins by introducing a graph structure that represents the topology of CDPRs, followed by network learning on the graph via message propagation block. The experimental results demonstrate that CafkNet can learn the inherent topological relationships of CDPRs and efficiently solve the FK issue. The results are validated across different types of CDPRs, including spatial CDPRs and planar CDPRs, as well as under configurations such as under-constrained, fully-constrained, and over-constrained CDPRs. The validation is conducted in both simulation environments and real-world scenarios as well.
This work focuses solely on the FK problem, which is known to be challenging in the kinematic modeling of CDPRs. So, in the future, we aim to apply the GNN to solve the IK problem and dynamic modeling issues.

### v4

@@ Abstract
Cable-driven parallel robots (CDPRs) have gained significant attention due to their promising advantages. When deploying CDPRs in practice, the kinematic modeling is a key question. Unlike serial robots, CDPRs have a simple inverse kinematics problem but a complex forward kinematics (FK) issue. So, the development of accurate and efficient FK solvers has been a prominent research focus in CDPR applications. By observing the topology within CDPRs, in this paper, we propose a graph-based representation to model CDPRs and introduce CafkNet, a fast and general FK solving method, leveraging Graph Neural Network (GNN) to learn the topological structure and yield the real FK solutions with superior generality, high accuracy, and low time cost. CafkNet is extensively tested on 3D and 2D CDPRs in different configurations, both in simulators and real scenarios. The results demonstrate its ability to learn CDPRs’ internal topology and accurately solve the FK problem. Then, the zero-shot generalization from one configuration to another is validated. Also, the sim2real gap can be bridged by CafkNet using both simulation and real-world data. To the best of our knowledge, it is the first study that employs the GNN to solve FK problem for CDPRs. Videos and codes are available at https://sites.google.com/view/cafknet/site.
@@ I Introduction
Cable-driven parallel robots (CDPR) is a kind of parallel mechanism that employs cables to replace rigid links [1, 2, 3, 4]. It has exhibited excellent scalability in terms of size, payload, and dynamics capacities.
The crucial step before deploying CDPRs is modeling, including kinematic modeling (KM) and dynamics modeling (DM). KM refers to the process of creating a mathematical equation of the end-effector movement, including its kinematic parameters such as position, orientation, and velocity. DM, on the other hand, is the process of creating a mathematical representation of how the end-effector will respond to external forces over time. For the KM of CDPR, the usual practice, like [5], is to first determine the configuration of the CDPR according to the application requirements, including the number of cables, positions of the pulley on the framework, the shape of the end-effector and the position on which the cable is connected. Then, according to the geometric relationship, the modeling of cable attachment on the end-effector and anchor points on the framework is formulated. After that, some downstream tasks can be completed, such as workspace determination [6], path planning [7], and control [8]. In CDPR’s KM, there are typically two types of calculations: forward kinematics (FK) and inverse kinematics (IK). FK involves determining the end-effector pose (i.e., position and orientation) based on cable lengths, while IK solves for the length of cables given the end-effector pose in reverse. In contrast to the traditional serial mechanical structure, the FK problem of CDPRs is more complex than its IK.
Therefore, numerous works focus on solving FK problems of CDPRs. [9] discusses the non-trivial computation of the FK of CDPRs, which are kinematically over-constrained mechanisms. In addition, [1] provides the overview of FK solvers based on successful application cases of CDPRs. Recent work [10] proposes a neural network (NN) approach for solving the FK of under-constrained CDPRs. In summary, solving the FK for CDPRs presents several challenges.
@@
Due to the nonlinearity and complexity of FK problems, FK solvers have a large computational burden and also suffer from the problem of no solution or multiple solutions. But in practical applications, people may focus on the real solution determination and not ask for all solutions, which, however, is the main purpose of existing FK solvers. Furthermore, configuration-dependent FK solvers may not be easily applied to the reconfiguration CDPR [11] or the case of adding or subtracting cables. However, according to our observations in [12], we notice that the change in CDPR’s configuration does not affect the topological architecture if its configuration can be represented as a graph. In addition, it is an indisputable fact that the Sim2Real gap is inevitable, despite significant efforts to reduce differences [13, 14, 15]. This is due to the difficulty of measuring and modeling systematic errors (e.g., cable attachment position errors) and random errors (e.g., cable sagging) in CDPR, some of which are even unknowable. One promising approach currently being explored is using real data to learn about these errors, as it is an effective way to bridge the gap.
In this paper, we propose a forward kinematics solving method based on graph neural Networks (GNN) for Cable-driven parallel robots, denoted by CafkNet. By utilizing the proposed graph representation for CDPR, CafkNet can rapidly learn its underlying topological structure. Due to the valid training data involved in GNN, the trained CafkNet can efficiently solve the real solutions to the FK problem, rather than finding all solutions. Thanks to the topological similarity of graphs, CafkNet trained from a specific configuration can be easily transferred to other configurations without fine-tuning. Additionally, CafkNet performs well in bridging the Sim2Real gap between vast easily accessible simulation data and limited expensive real-world data. We have conducted extensive simulations and real-world experiments to demonstrate the modeling capabilities of CafkNet.
@@
  • We propose the CafkNet, a fast and general method to yield real solutions to the FK problem of CDPRs with any configurations via GNN.
@@
The overview of CafkNet is exhibited in Fig. 1. The rest of the paper is organized as follows. Sec. II reviews related studies about KM of CDPR and GNN applications in robotics. Sec. III introduces the proposed graph representation of CDPR and the architecture of CafkNet. Then, Sec. IV demonstrates the experiment details followed by modeling results of CafkNet in Sec. V. Finally, the conclusion and outlook are presented in Sec. VI.
Figure 1: Overview of CafkNet. (a) The geometric model of a spatial CDPR with $m$ cables. (b) The proposed graph presentation for the CDPR (see Sec. III-C). (c) The top-down computation scheme in each message propagation block. The cable nodes collect data from the world node and are eventually aggregated by the body node (see Sec. III-D).
@@ II-A Kinematic Modeling of CDPR
KM is a fundamental problem in CDPR’s studies and also the first step in each CDPR’s application, as revealed in [1, 2, 3]. In CDPR, the IK problem, i.e., the rope length determination given the end-effector pose, is easy to solve, relative to the FK issue. Due to the nonlinearities in FK, there are three potential situations when solving FK [16], i.e., no solution, one solution, and multiple solutions. In addition, [9] introduces an FK solver based on energy minimization for a numerical optimization technique. Also, [17] takes into account the nonlinearities of cables and proposes a method for solving the FK of CDPRs with sagging cables. However, these methods are mainly used to model CDPR more finely under their respective configurations, so as to complete the optimization of FK. More generally, some studies [18, 17] provide the FK solver using graph and graph theory to model the CDPR. Recent works employ diverse machine learning methods trying to solve the system modeling problem, such as neural networks [19, 10], deep learning [20], etc. Although their results validate the feasibility of neural networks on FK, they can not be directly applied to new configurations and need to be retrained when changing robot configurations. In this work, we combine the graph and neural network and utilize GNN to represent and model the CDPR kinematics. Benefitting from the graph similarity, the proposed CafkNet could transfer its learned topological structure of CDPR to other configurations without retraining. This can not be done by NN-based FK solvers.
@@ II-B Graph Neural Networks in Robotics
Recently, data-driven methods have emerged as an effective approach to approximate robotics modeling. Given the fact that the robot’s structure can be easily represented as a graph, where joints are denoted as nodes and links are denoted as edges, GNNs have been increasingly applied in robotics sensing [21, 22], modeling [23, 24, 25, 26], and control [27, 28, 29, 30].  [21, 22] propose to use GNN to encode the distributed tactile sensing information for downstream tasks. Some works employ GNN to model system dynamics, where graphs represent the rigid body [23, 24] and soft cloth [25]. [26] aims to find the representation of robot data, including kinematic data and motion data using the GNN message computation scheme. NerveNet [27] seeks to train a generalized control policy that can transferred to other robot configurations.  [28, 29, 30] propose a similar idea to train a structure-agnostic locomotion policy with a novel message-passing scheme. In this paper, we employ the graph to represent the structure of CDPRs and train an FK model based on GNN. Since the training data only contains valid data about cable length and pose, the trained CafkNet can directly find the real solutions to the FK problems rather than struggling in all solutions determination. To the best of our knowledge, it is the first study to employ the GNN for FK problem of CDPRs.
III Methodology
@@ III-A Geometric Model of CDPR
Here we consider the tightrope in CDPRs so that the cable model can be represented as the vector. A 6-degree-of-freedom (DoF) spatial CDPR is illustrated in Fig. 1-(a), where two endpoints of the $i$-th cable ($i=1,2,\cdots,m$) are denoted by $A_{i}$ and $B_{i}$. Given the geometric relationship, the cable vector $\mathbf{l}_{i}$ can be determined as follows.
@@ III-B Kinematics of CDPR
III-C Graph Representation of CDPR
@@
III-D GNN-Empowered Forward Kinematic Modeling
@@
III-E Transfer Learning across Configurations
CafkNet employs parameter sharing among identical node relationships, enhancing its ability to perform structural transfer tasks effectively. Specifically, there are only two parameterized MLPs in our graph for message propagation, including $\gamma_{wc}$ for the relationship between world node and cable node, as well as $\gamma_{cb}$ for the relationship between cable node and body node. The parameters of MLPs are independent of cable number. As a result, our approach facilitates the application of the kinematic model to various CDPR configurations, significantly improving generalization capabilities across different cable arrangements. Extensive experiments in Sec. V verify the transferability of our CafkNet.
IV Experiments
IV-A Dataset
@@ IV-C Implementation Details
V Results and Discussion
V-A FK Modeling for CDPR
@@
V-B Transfer Learning in CDPR
@@ V-C Bridging Sim2Real Gap
V-C1 Simulated Data with Artificial Noises
@@
V-C2 Real Experimental Data
@@
VI Conclusion
This paper presents the CafkNet, a GNN-empowered FK-solving method for CDPRs with superior generality, high accuracy, and low time cost. It is the first work to utilize GNN to handle the FK issue of CDPRs. The study begins by introducing a graph structure that represents the topology of CDPRs, followed by network learning on the graph via message propagation block. The experimental results demonstrate that CafkNet can learn the inherent topological relationships of CDPRs and efficiently find real solutions to the FK problem. The results are validated across different types of CDPRs, including spatial CDPRs and planar CDPRs, as well as under configurations such as under-constrained, fully-constrained, and over-constrained CDPRs. The validation is conducted in both simulation environments and real-world scenarios. This work only considers the straight cables, so the appropriate graph representation for sagging cables in FK problem may be a future work.